# High Frequency Quoting Under Liquidity Constraints

### v1

@@ 9 Forecasting Execution Risk with Multivariate Hawkes Processes
The preceding section introduced the structure of tick-level events segmented by order book side and event type. We now present a predictive framework that leverages these microstructural observations to forecast execution risk across candidate reference legs. Specifically, we construct forward-looking arrival ratios using multivariate Hawkes processes and use them to infer a reference leg decision. This model-based strategy complements the reactive rule-based approaches presented earlier and provides a direct forecasting proxy for the oracle decision $\chi^{m}(w)$ introduced in Problem 1.2.
9.1 Forecasting Objective and Model Structure
@@
Let =$\mathcal{E}=\{T,C,PDM\}$ denote the set of reference-impacting event types: trades, cancellations, and downward price modifications. Let $Y=\{b,a\}$ represent the bid and ask sides of the order book, and $X\in\{c,n\}$ index the current and next-month contracts. Define the joint index set =×$\mathcal{I}=\mathcal{E}\times Y$.
@@
To maintain consistency with the decision framework introduced in Problem 1.2, we denote this forecasting-driven rule as $\mathcal{Q}^{\mathcal{H}}$, with its output defined as:
$\chi^{\mathcal{H}}(w):=\tilde{\chi}_{(\tau,\tau+\xi]}=\mathbb{I}\left\{\rho^{c,a}_{(w)}\geq\rho^{n,b}_{(w)}\right\},\quad\text{where }w=(\tau,\tau+\xi]$
Here, $\mathcal{Q}^{\mathcal{H}}$ maps estimated arrival ratios to a binary reference leg decision over window $w\in\mathcal{W}$. This rule completes the triad of decision functions $\{\mathcal{Q}^{\mathcal{T}},\mathcal{Q}^{\mathcal{B}},\mathcal{Q}^{\mathcal{H}}\}$, all evaluated against the market oracle $\chi^{m}(w)$ using the agreement score defined in Problem 1.2.
Formal Definition: Hawkes-Derived Feature Mapping. For each interval $w=[w_{1},w_{2})\in\mathcal{W}$, let $\mathcal{G}^{\mathcal{H}}(w)\in\mathbb{R}^{2}$ denote the filtered representation obtained by simulating a fitted Hawkes model on recent tick data to estimate directional arrival ratios. Define: $\tilde{N}^{X,Y}_{(w)}\quad\text{Simulated total Hawkes event counts on }w$ $\tilde{N}^{X,Y,r}_{(w)}\quad\text{Simulated events attributed to leg }r\in\{c,n\}$ $\rho^{X,Y}_{(w)}:=\frac{\tilde{N}^{X,Y,r}_{(w)}}{\tilde{N}^{X,Y}_{(w)}}$ $\mathcal{G}^{\mathcal{H}}(w):=\left(\rho^{c,a}_{(w)},\;\rho^{n,b}_{(w)}\right)$ $\mathcal{Q}^{\mathcal{H}}(w):=\arg\min_{r\in\{c,n\}}\rho^{r,s}(w)$ $\chi^{\mathcal{H}}(w):=\mathbb{I}\left\{\mathcal{Q}^{\mathcal{H}}(w)=n\right\}$ where $\rho^{r,s}(w)$ represents the simulated relative arrival intensity for quoting side $s$ of leg $r$, and $n$ refers to the next-month contract.
The above formalization defines the Hawkes-based decision rule $\chi^{\mathcal{H}}(w)$ as a function of simulated forward event counts derived from fitted multivariate Hawkes models. By comparing the relative frequency of reference-price-impacting events across book sides, this approach translates microstructural event flow forecasts into a structured quoting decision. This provides a direct, data-driven proxy for short-term execution stability under each reference leg and can be evaluated against the market-optimal decision $\chi^{m}(w)$.
Modeling Remark This decision rule $\chi^{\mathcal{H}}(w)$ serves as a forecasting-based analog of the oracle decision $\chi^{m}(w)$. It selects the contract whose book side is expected to exhibit fewer reference-price-impacting events, thus offering more stable execution.
9.2 Multivariate Hawkes Process for Event Forecasting
@@
9.3 Estimation and Simulation Procedure
@@
Non-parametric kernels, by contrast, do not assume a fixed functional form. Instead, we consider two classes of non-parametric estimators: (i) the expectation-maximization (EM) method, which iteratively estimates a discretized kernel by maximizing a latent-variable likelihood, and (ii) the conditional law approach, which reconstructs the kernel by inverting the conditional intensity relation from observed event times. These methods provide flexible approximations of $\phi_{ij}(t)$, allowing for data-driven recovery of potentially complex excitation structures.
@@
Hawkes kernels are estimated via maximum likelihood using the method of [20]. For parametric kernels, we use exponential and sum-of-exponentials forms. Non-parametric estimation follows [21, 22]. Simulated event sequences are generated using Ogata’s thinning method [23].
Tick history $(\tau-h,\tau]$Filter: =$\mathcal{E}=\{T,C,PDM\}$Estimate kernel $\hat{\Phi}^{X,Y,r}$imulate $\tilde{N}^{X,Y,r}$o filtering: Use all ticksEstimate kernel $\hat{\Phi}^{X,Y}$imulate $\tilde{N}^{X,Y}$ompute arrival ratio $\rho^{X,Y}_{(\tau,\tau+\xi]}$eference leg decision $\chi^{\mathcal{H}}(w)$
@@
9.4 Forecast Evaluation and Model Selection
@@
The best-performing kernel $\hat{\Phi}^{*}$ is used for all subsequent simulations. The resulting forecasting-based decision rule $\chi^{\mathcal{H}}(w)$ is benchmarked against the oracle decision $\chi^{m}(w)$ using the agreement score defined in Problem 1.2.
@@
Summary: Integration into Evaluation Framework The Hawkes-based arrival ratio forecast yields a reference leg decision that anticipates short-term slippage risk. This decision is validated by comparing its alignment with the oracle decision across evaluation windows, allowing rigorous benchmarking against book- and tick-based quoting rules.
10 Reference Leg Selection via Order Book Liquidity Profiles
The preceding section developed a reference leg selection rule based on intensity forecasts derived from tick history using multivariate Hawkes processes. In contrast, we now consider a structurally simpler approach that relies entirely on the instantaneous state of the order book—specifically, the slope and depth of price levels on each side. While the Hawkes-based method models temporal dependencies in event arrivals, the order-book-state-based method assumes that key execution-relevant information is already embedded in the current liquidity profile. This contrast implicitly reflects two different informational assumptions: one emphasizes historical flow; the other emphasizes present structure. We make no claims about market efficiency, but use this contrast to motivate an alternative, real-time computable quoting rule based on liquidity discontinuities.
While recent models of limit order book dynamics—particularly those based on multivariate Hawkes processes—offer detailed descriptions of market microstructure, their practical adoption in high-frequency trading (HFT) remains limited. This gap stems primarily from the computational complexity inherent in estimating and updating such models in real time. Studies such as Chavez-Demoulin and McGill [25] emphasize the intensive nature of maximum likelihood estimation for Hawkes processes, which can render them unsuitable for latency-sensitive HFT environments. Real-time simulation of event intensities, which depend on the full event history, further adds to the processing burden, especially in multivariate settings where many event types and cross-excitation terms are involved.
The trade-off between model expressiveness and tractability becomes especially pronounced in quoting strategies that require continuous adaptation. Consequently, practical HFT systems often rely on simplified heuristics or lightweight statistical indicators that can be computed at sub-millisecond timescales. This motivates the development of forecasting frameworks that retain the interpretability and structure of LOB-aware models—such as those based on Hawkes processes—while offering tractability and responsiveness suitable for real-time quoting systems [19]. However, a key limitation of this framework is its exclusion of the absolute price levels present in the LOB. While the Hawkes model effectively represents the timing and interaction of events, it treats the price itself as exogenous and does not integrate its magnitude into the structure of the model.
This modeling lacuna has important implications when the objective is to identify the more stable reference contract for execution. In particular, discontinuities in the price levels across the order book can lead to non-trivial slippage, especially when order sizes exceed the available liquidity at the best quotes. For example, in Table 1, the bid side of the NIFTY February 2022 futures contract shows a significant drop from INR 17450 at price level 2 to INR 17401.1 at price level 3, on the bid side,—a INR 48.9 decline. A market order large enough to consume the top two bid levels would thus incur a sharp execution cost due to this gap.
The magnitude of such price-level drops is crucial for modeling expected slippage and, by extension, for determining the optimal reference contract under liquidity constraints. The current Hawkes-based framework, however, is not equipped to capture this dimension. This limitation suggests the need for an augmented model that accounts for price slope and liquidity profile.
@@ 10 Reference Leg Selection via Order Book Liquidity Profiles
The proposed Composite Liquidity Factor (CLF) builds on established work in empirical market microstructure that connects order book shape to execution cost. Hasbrouck’s concept of the log quote slope—defined as the rate of price change per unit of depth—serves as a key antecedent [26]. Figure 3 illustrates this slope as the gradient of the line connecting logarithmic quantity and price at the top levels of the bid and ask. As more quantity accumulates near the mid-price, this slope flattens, indicating increased liquidity. Bouchaud et al. [27] and Cont and de Larrard [28] further emphasize how convexity and depth regularities shape price impact and short-term stability. Glosten’s decomposition of spreads into permanent and transitory components [29] also links slope to the informational cost of trading. The CLF adapts these ideas into a direction-sensitive liquidity metric, tailored to identify the more execution-stable reference leg in rollover spreads.
@@
Figure 3: Illustration of the log quote slope: the slope of the dashed line connecting $(-\log(q^{b,i}),p^{b,i})$ and $(\log(q^{a,i}),p^{a,i})$. Adapted from Hasbrouck and Seppi (2001, Fig. 2).
The log-quote slope is a well-established measure of liquidity that combines information from the bid-ask spread and the associated quote depth. However, for the purpose of optimally selecting a reference contract in our setting, we are particularly interested in modeling liquidity asymmetries—specifically, discontinuities on a single side of the order book. The relevant side (bid or ask) depends on whether we are initiating a buy or sell spread.
In this paper, we consider the case of futures rollovers as our motivating example. A typical futures rollover strategy involves buying the spread, which means selling the current month’s futures contract and simultaneously buying the next month’s contract. To execute this rollover with minimal slippage, it is crucial to evaluate which side—either the bid side of the current month contract or the ask side of the next month contract—offers better instantaneous liquidity for the required quantity.
To this end, we introduce a modified formulation of the log-quote slope tailored to our needs. We refer to this new measure as the Composite Liquidity Factor (CLF). This formulation retains the essence of the original slope measure but adapts the notation to remain consistent with the rest of the paper.
When using the current month futures contract as the reference leg, the CLF is defined as:
$CLF^{b,i}_{c}=\frac{\log(p^{b,1}_{c}/p^{b,i+1}_{c})}{\log(\sum_{j=1}^{i}q^{b,j}_{c})}\quad\forall\ p^{b,i+1}_{c}>p^{b,i}_{c},i+1\leq\nu$ (22)
Conversely, when using the next month futures contract as the reference leg, the CLF is defined as:
$CLF^{a,i}_{n}=\frac{\log(p^{a,i+1}_{n}/p^{a,1}_{n})}{\log(\sum_{j=1}^{i}q^{a,j}_{n})}\quad\forall\ p^{a,i}_{n}>p^{a,i+1}_{n},i+1\leq\nu$ (23)
@@
Here, $p^{Y,i}_{X}$ represents the price at level $i$ of the relevant side of the order book, and $q^{Y,j}_{X}$ denotes the aggregated quantity to level $i$. The numerator captures the logarithmic price drop (or rise), while the denominator aggregates the logarithmic depth over the first $i$ levels. This construction provides a flexible and direction-sensitive indicator of slippage risk, suited to the decision of reference contract selection in spread-based strategies. Unlike the Hawkes-based approach, the CLF-based rule requires no model fitting or simulation. All quantities used in the decision rule are computed directly from the current limit order book snapshot, making the method suitable for real-time application in latency-sensitive environments.
@@
We now define the CLF-based quoting rule formally, using the same decision function structure introduced in Problem 1.2.
@@
Formal Definition: CLF-Based Reference Leg Selection Rule. For each evaluation window $w=[w_{1},w_{2})\in\mathcal{W}$, let $\mathcal{G}^{\mathcal{B}}(w)\in\mathbb{R}^{2}$ denote the book-based filter output containing directional liquidity values for the two contract legs. Define: $\mathcal{G}^{\mathcal{B}}(w):=\left(CLF_{c}(w),\;CLF_{n}(w)\right)$ $\mathcal{Q}^{\mathcal{B}}(w):=\arg\min_{r\in\{c,n\}}CLF_{r}(w)$ $\chi^{\mathcal{B}}(w):=\mathbb{I}\left\{\mathcal{Q}^{\mathcal{B}}(w)=n\right\}$ where $CLF_{r}(w)$ denotes the composite liquidity factor for leg $r\in\{c,n\}$, and $n$ refers to the next-month contract.
@@
This decision rule operates entirely on the book-based filtered representation $\mathcal{G}^{\mathcal{B}}(w)$, requiring no historical data, simulation, or model fitting. It is fully deterministic and computed from observable execution risk at the quoting time $w_{1}$. By capturing asymmetries in directional liquidity across contract legs, $\mathcal{G}^{\mathcal{B}}(w)$ provides a structurally simpler, real-time quoting signal that complements event-driven methods.

### v2

@@ 3.4 Forecasting Execution Risk with Multivariate Hawkes Processes
Building on the event taxonomy above, we introduce a predictive framework that forecasts short-horizon execution risk across candidate reference legs. We estimate forward-looking arrival ratios with multivariate Hawkes processes and map them to a reference-leg decision. This model-based rule complements the reactive rules presented earlier and serves as a forecasting proxy for the oracle decision $\chi^{m}(w)$ defined in the Problem Definition (Section 2). The next subsection specifies the event sets, sides, and windows used to form these arrival-ratio forecasts.
Model Structure
@@
Let $\mathcal{E}=\{T,C,PDM\}$ denote the set of reference-impacting event types: $\mathcal{T}$ for trades, $\mathcal{C}$ for cancellations, and $\mathcal{PDM}$ for downward price modifications. Let $Y=\{b,a\}$ represent the bid and ask sides of the limit order book, and $X\in\{c,n\}$ index the current and next-month contracts. Define the joint index set $\mathcal{I}=\mathcal{E}\times Y$.
@@
To maintain consistency with the decision framework introduced in Section 2, we denote this forecasting-driven rule as $\mathcal{Q}^{\mathcal{H}}$, with its output defined as:
$\chi^{\mathcal{H}}(w):=\tilde{\chi}_{(\tau,\tau+\xi]}=\mathbb{I}\left\{\rho^{c,a}_{(w)}\geq\rho^{n,b}_{(w)}\right\},$
$\quad\text{where }w=(\tau,\tau+\xi]$
Here, $\mathcal{Q}^{\mathcal{H}}$ maps estimated arrival ratios to a binary reference leg decision over window $w\in\mathcal{W}$.
[…]
The above formalization defines the Hawkes-based decision rule $\chi^{\mathcal{H}}(w)$ as a function of simulated forward event counts derived from fitted multivariate Hawkes models. By comparing the relative frequency of reference-price-impacting events across LOB sides, this approach translates microstructural event flow forecasts into a structured quoting decision. This provides a direct, data-driven proxy for short-term execution stability under each reference leg and can be evaluated against the market-optimal decision $\chi^{m}(w)$.
Modeling Remark This decision rule $\chi^{\mathcal{H}}(w)$ serves as a forecasting-based analog of the oracle decision $\chi^{m}(w)$. It selects the contract whose LOB side is expected to exhibit fewer reference-price-impacting events, thus offering more stable execution.
Multivariate Hawkes Process for Event Forecasting
@@
Estimation and Simulation Procedure
@@
Non-parametric kernels, by contrast, do not assume a fixed functional form. Instead, we consider two classes of non-parametric estimators: (i) the expectation-maximization (EM) method, which iteratively estimates a discretised kernel by maximizing a latent-variable likelihood, and (ii) the conditional law approach, which reconstructs the kernel by inverting the conditional intensity relation from observed event times. These methods provide flexible approximations of $\phi_{ij}(t)$, allowing for data-driven recovery of potentially complex excitation structures.
@@
Hawkes kernels are estimated via maximum likelihood using the method presented in [40]. For parametric kernels, we use exponential and sum-of-exponentials forms. Non-parametric estimation follows [3, 33]. Simulated event sequences are generated using Ogata’s thinning method [38].
Tick history $(\tau-h,\tau]$Filter: $\mathcal{E}=\{T,C,PDM\}$Estimate kernel $\hat{\Phi}^{X,Y,r}$Simulate $\tilde{N}^{X,Y,r}$No filtering: Use all ticksEstimate kernel $\hat{\Phi}^{X,Y}$Simulate $\tilde{N}^{X,Y}$Compute arrival ratio $\rho^{X,Y}_{(\tau,\tau+\xi]}$Reference leg decision $\chi^{\mathcal{H}}(w)$
@@
Forecast Evaluation and Model Selection
@@
The best-performing kernel $\hat{\Phi}^{*}$ is used for all subsequent simulations. The resulting forecasting-based decision rule $\chi^{\mathcal{H}}(w)$ is benchmarked against the oracle decision $\chi^{m}(w)$ using the agreement score defined in Section 2.
@@
Summary: Integration into Evaluation Framework The Hawkes-based arrival ratio forecast yields a reference leg decision that anticipates short-term slippage. This decision is validated by comparing its alignment with the oracle decision across evaluation windows, allowing rigorous benchmarking against LOB- and tick-based quoting rules.
3.5 LOB-based quoting rule.
Similarly, for each $w\in\mathcal{W}$, we define a LOB-based quoting rule that uses the limit order book snapshot observed at time $w_{1}$ to evaluate expected slippage:
with associated decision function:
These rules instantiate the framework in Section 2 using real-time market observables. The tick-based rule aggregates historical event data, and the LOB-based rule extracts instantaneous liquidity conditions. Both decisions use only information timestamped strictly before $w_{1}$. The realized outcomes on $w_{1},w_{2})$ are used solely for evaluation. When the two leg scores are exactly equal, we break ties in favor of $c$ for determinism and comparability with the oracle.
3.6 Reference Leg Selection via LOB Liquidity Profiles
Building on the previous section’s reference-leg selection rule derived from multivariate Hawkes intensity forecasts, we now introduce a LOB-state alternative that relies solely on the instantaneous limit order book configuration. Concretely, we aggregate level-wise depth with cross-level price gaps to form a composite liquidity profile (defined below). Whereas the Hawkes arrival ratio encodes temporal dependencies in event flow, this LOB-state construction encodes contemporaneous structure. We use this contrast to motivate a real-time, reference-leg selection rule based on liquidity discontinuities, with the empirical role of LOB-state information assessed against our evaluation oracle (deviation from the target spread).
Recent models of limit order book dynamics, including multivariate Hawkes formulations, offer detailed descriptions of event arrivals and state dependence [15, 41, 36]. In latency-sensitive HFT deployments, real-time estimation and updating can be computationally demanding, particularly as the number of event types and cross-excitation terms grows [15]. This motivates a complementary rule that is computable at decision times from LOB-state features alone.
Motivated by the computational considerations in the preceding paragraph, we make explicit a practical trade-off between model expressiveness (rich event-history dependence) and runtime tractability (updates computable within each decision interval). In latency-sensitive deployments, low-latency decision rules with tractable state updates are commonly employed [14, 2]. Full re-estimation or online updating of high-dimensional event-history models can be computationally demanding [15]. We therefore consider a complementary LOB-state framework that preserves LOB-aware structure while remaining tractable at runtime. As defined in Section 3 and detailed in Section 3.4, the Hawkes arrival-ratio construction models the timing of events from history and treats the absolute price level as exogenous.
Grounding the tractability discussion above in a concrete LOB-state effect, cross-level price gaps can produce materially higher execution cost when available depth at the best quotes is insufficient [17, 10]. In Table 1, the bid ladder exhibits a drop from —17450 at level 2 to —17401.1 at level 3, a —48.9 gap. An order that consumes the top two bid levels and then executes at level 3 would incur approximately —48.9 additional cost per contract due to this discontinuity. This motivates encoding price levels and cross-level gaps directly in our LOB-state measure for reference-leg selection, formalized in Section 3.6.
Extending the example above, the magnitude of cross-level price drops directly determines execution cost when top-of-book depth is insufficient, as illustrated in Table 1. Because our Hawkes arrival-ratio in Section 3.4 treats the absolute price level as exogenous, it is not directly sensitive to these discontinuities. We therefore introduce a complementary LOB-state measure that encodes price levels, cross-level price gaps, and level-wise depth, and we assess its effect on reference-leg selection using the oracle benchmark defined in Section 3.6.
@@ 3.6 Reference Leg Selection via LOB Liquidity Profiles
Building on the need to encode price levels, cross-level gaps, and depth stated above, the Composite Liquidity Factor (CLF) draws on established links between limit order book shape and execution cost. Hasbrouck’s log quote slope (the rate of price change per unit depth) provides a core antecedent [28]. Figure 3 depicts this slope as the gradient of the line connecting logarithmic quantity and price at the top levels; a greater quantity near the mid-price corresponds to a flatter slope and higher immediate liquidity. Prior work shows how shape and depth regularities affect short-horizon price impact and stability [10, 17], while Glosten’s decomposition of spreads links such shape effects to informational trading costs [21]. We adapt these ideas into a direction-sensitive LOB-state measure (CLF) for reference-leg selection.
@@
Figure 3: Illustration of the log quote slope: the slope of the dashed line connecting $(-\log(q^{b,i}),p^{b,i})$ and $(\log(q^{a,i}),p^{a,i})$. Adapted from [28], Figure 2.
Although the log-quote slope combines spread and depth across both sides of the LOB [28], reference-leg selection in our setting is side-conditional: we require a measure that emphasizes one-sided liquidity asymmetries across price gaps and level-wise depth on the side that will be consumed. The relevant side (bid or ask) is determined by whether the spread order is a buy or a sell. We therefore construct CLF features to encode these one-sided discontinuities (Section 3.6) and evaluate their contribution using deviation from the target spread.
To operationalise the side-conditional requirement above, we introduce a LOB-state measure that adapts the log-quote slope to one-sided execution. We refer to this measure as the Composite Liquidity Factor (CLF). CLF preserves the slope intuition while explicitly encoding price levels, cross-level price gaps, and level-wise depth on the side to be consumed; formal definitions follow in Section 3.6.
Let $p^{Y,i}_{X}$ denote the price at level $i$ on side $Y\in\{b,a\}$ (bid or ask) for contract $X\in\{c,n\}$ (current or next month), and let $q^{Y,j}_{X}$ denote the available quantity at level $j$. With the current-month contract used as the reference leg, we define a bid-side CLF at depth $i$ that contrasts the cross-level price gap with cumulative available depth on the consumed side:
$CLF^{b,i}_{c}\;=\;\frac{\log\!\big(p^{b,1}_{c}/p^{b,i+1}_{c}\big)}{\log\!\big(\sum_{j=1}^{i}q^{b,j}_{c}\big)},\qquad i\in\{1,\dots,\nu\!-\!1\},\;\;p^{b,i+1}_{c}<p^{b,i}_{c}.$ (22)
Here, the numerator encodes the price drop from the best bid to level i+1i\!+\!1, and the denominator encodes cumulative bid-side depth through level $i$.
Conversely, for the next-month contract used as the reference leg, we define an ask-side CLF at depth $i$ that contrasts the cross-level price gap with cumulative available depth on the consumed side:
$CLF^{a,i}_{n}\;=\;\frac{\log\!\big(p^{a,i+1}_{n}/p^{a,1}_{n}\big)}{\log\!\big(\sum_{j=1}^{i}q^{a,j}_{n}\big)},\qquad i\in\{1,\dots,\nu\!-\!1\},\;\;p^{a,i+1}_{n}>p^{a,i}_{n}.$ (23)
Here, the numerator encodes the price increase from the best ask to level i+1i\!+\!1, and the denominator encodes cumulative ask-side depth through level $i$.
@@
This construction provides a flexible and direction-sensitive indicator of slippage risk, suited to the decision of reference contract selection in spread-based strategies. Unlike the Hawkes-based approach, the CLF-based rule requires no model fitting or simulation. All quantities used in the decision rule are computed directly from the current limit order book snapshot, making the method suitable for real-time application in latency-sensitive environments.
@@
We now define the CLF-based quoting rule formally, using the same decision function structure introduced in Section 2.
@@
Definition: CLF-Based Reference Leg Selection Rule. For each evaluation window $w=[w_{1},w_{2})\in\mathcal{W}$, let $\mathcal{G}^{\mathcal{B}}(w)\in\mathbb{R}^{2}$ denote the LOB-based filter output containing directional liquidity values for the two contract legs. Define: $\mathcal{G}^{\mathcal{B}}(w):=\left(CLF_{c}(w),\;CLF_{n}(w)\right)$ $\mathcal{Q}^{\mathcal{B}}(w):=\arg\min_{r\in\{c,n\}}CLF_{r}(w)$ $\chi^{\mathcal{B}}(w):=\mathbb{I}\left\{\mathcal{Q}^{\mathcal{B}}(w)=n\right\}$ where $CLF_{r}(w)$ denotes the composite liquidity factor for leg $r\in\{c,n\}$, and $n$ refers to the next-month contract.
@@
This decision rule operates entirely on the LOB-based filtered representation $\mathcal{G}^{\mathcal{B}}(w)$, requiring no historical data, simulation, or model fitting. It is fully deterministic and computed from observable execution risk at the quoting time $w_{1}$. By capturing asymmetries in directional liquidity across contract legs, $\mathcal{G}^{\mathcal{B}}(w)$ provides a structurally simpler, real-time quoting signal that complements event-driven methods.